Two masses, one of 2.5 kg and one of 1.5 kg are connected by a light cord passing over a pulley of 150 mm diameter and mass 1 kg, which may be considered as a uniform disc Calculate the acceleration of the masses and the tension in the cord when the system is released from rest.


1
Expert's answer
2021-07-12T12:16:01-0400
"(m+M)a+\frac{I\alpha}{r}=(M-m)g\\
(m+M+0.5m_d)a=(M-m)g\\
(2.5+1.5+0.5)a=(2.5-1.5)9.8\\a=2.2\frac{m}{s^2}\\T=2.5(9.8-2.2)=19\ N"